OpenStudy (anonymous):

Is the vector (b1) (1) (1 ) (1) (b2)=(0)+c1(0 )+c2(0) (b3) (0) (-1) (1) a subspace? The video for the Recitation on Vector Subspaces says it is, but somehow it does not seem to contain the zero vector.

OpenStudy (anonymous):

try c1=-1/2 and c2=-1/2

OpenStudy (anonymous):

Thank you! I get it now! :)